A body is projected at anangke of 30degrees to the horizontal with a velocity of 39.2 m/s The ratio of its vertical components of velocity after 1s,3s and 4s is a)1:-1:-2 b)-1:2:-3 c)1:2:3 d)1:2:1

Grade:12

1 Answers

Adarsh
6 years ago
Dear, We can calculate it by finding vertical component of velocity after 1,3,4 second .Vertical component of velocity= 39.2 .sin30°=19.6 m/s.Velocity after 1 second = 19.6 - 9.8 =9.8 Velocity after 3 second = 19.6 - 9.8 . 3= -9.8Velocity after 4 second= 19.6 - 9.8 . 4 = -19.6 Now ratio of velocity = 1: -1 : -2 Hence (a) is the correct option